OpenStudy (buttermequeasy):

Choose the equation below that represents the line passing through the point (1, -4) with a slope of 1/2. y - 4 = 1/2(x - 1) y + 4 = 2(x - 1) y + 4 = 1/2(x - 1) y - 4 = 2(x + 1)

OpenStudy (buttermequeasy):

help please!

OpenStudy (mysesshou):

(1, -4) = (x1,y1) 1/2 = m. equation of the format: (y-y1) = m* (x-x1)